OpenStudy (anonymous):

Can someone check my work on this? A truck carrying a 38 kg crate accelerates uniformly from rest to 87.5 km/hr (<-- which I converted to 24.31 m/s) in 12.1 s. Calculate the work done on the crate by the truck. force = (38)(9.8) = 372.4 distance = 24.31*21 = 510.51 force * distance = 190113.9 = work (What units would this be?)

OpenStudy (anonymous):

Force = m x a where m is mass of truck, and a is acceleration of truck a = (vt -vo)/t where vt is the final speed, vo is the initial speed, and t is time., a = (24.31 m/s - 0)/ 12.1 s a = 2 m/s^2 So Force = m x a Force = 38 kg x 2 m/s^2 Force = 76 N Finally Work = F x s Work = 76 N x (vo t + 1/2 at^2) m Work = 76 N x (0+ (0.5 x 2 x 12.1^2)) m Work = 76 N x 146.41 m Work = 11127.16 Joule

OpenStudy (anonymous):

@sponkers_w ???

OpenStudy (anonymous):

ohhhhh okay. I guess I need to practice the acceleration bit more.

OpenStudy (anonymous):

@sponkers_w check this out http://www.physicsclassroom.com/Class/1DKin/U1L1e.cfm maybe thats useful

OpenStudy (anonymous):

@sponkers_w you forgot to give me a medal :)